# Dielectric waveguide setup tested with a superconducting millimeter-wave Fabry-Pérot interferometer at milli-Kelvin temperatures

### v1

@@ Abstract
We proposed and tested a cryogenic setup comprising dielectric waveguides for mm-wave frequencies in the range of GHz and temperatures down to 10 mK. The targeted applications are quantum technologies at millimeter-wave frequencies, which require measurements at low photon numbers and noise. We show that the high density polyethylene waveguides combine a frequency independent low photon loss with a very low heat conductance. Black high density polyethylene shows a higher attenuation, which is useful to block thermal photons in a cryogenic environment. The dielectric waveguides are thermally anchored and attenuated at several stages of the cryostat. They are individually protected by additional metallic shields to suppress mutual cross-talk and external interference. We have measured a Fabry-Pérot cavity with superconducting mirrors at 10 mK and found out that the quality of a signal transmitted through the dielectric waveguides is sufficient to measure resonator quality factors over one million at 110 GHz.
@@ I Introduction
Millimeter wave measurements at ultra low temperatures are widely unexplored due to several technical difficulties including, for instance, a challenging signal path thermalization and isolation. The tremendous progress in quantum computing at microwave frequencies seen over the last two decades did not so far approach the mm-wave range. The mm-wave quantum measurements would profit from an order of magnitude larger photon energy, wider bandwidth, enhanced resilience against thermal fluctuations, and may speed up qubit manipulation.
The traditionally used metallic hollow waveguides are very good conductors of mm-wave signals but are not flexible and cause an intolerable heat load on the cryogenic temperature stages. Also, the transmission of thermal radiation from elevated temperature stages can be problematic due to opening path to excess flux of infrared photons which heat up the low-temperature stages. Here, we propose an alternative approach to 100 GHz range quantum measurements by using dielectric waveguides (DWGs) to transmit mm-wave signals from room to milli-Kelvin temperatures and back. DWGs offer several advantages: An ultra low heat conductance, very good transmission properties, and can be easily thermalized. Furthermore, DWGs exhibit very good mechanical flexibility and offer several options to handle the attenuation and suppression of infrared radiation. Although it was long known that a dielectric strip can carry mm-waves [1], practical investigations motivated by advances in information technology were carried out only about one decade ago [2] and are still an attractive area of developments [3]. Very recent usage of DWGs for cryogenic applications employed frequencies up to 300 GHz [4]. There, a circuit board was used to bridge the temperature stages of a cryostat to a base temperature of 4.2 K.
In this paper, we describe a cryogenic setup including DWG signal lines with a high signal attenuation at the input and a low signal attenuation at the output lines. We also report on testing the setup with a superconducting Fabry-Pérot interferometer and observation of high-Q resonances in the W-band (75-110 GHz).
@@ II Ultra-low temperature millimeter-wave setup
The developed cryogenic mm-wave setup consists of four main parts: A commercial pulse tube pre-cooled dilution refrigerator (Blufors LD250) with a base temperature of 10 mK, a mm-wave vacuum-transition, black and transparent DWGs bridging the temperature stages and a low noise amplifier with an isolator located at the 2.7 K stage. Pre-characterization measurements are also carried out in a home-made dry refrigerator with a base temperature of 2.4 K (TransMIT 2-stage pulse-tube type PtQUBE). The Fabry-Pérot cavity, discussed in the second part of the paper, is directly connected to the DWGs, see Fig. 1 for a schematic representation. All presented measurements are carried out at room temperature using a Rhode & Schwarz ZNA67 vector network analyzer (VNA) with external W-band extenders (ZC110).
@@ II.1 Vacuum transition
We use one standard KF40 flange on the top of the fridge to add a vacuum tight box equipped with four copper WR10 waveguides. Each feed-through of the flange is closed by an o-ring sealed stainless steel metal plate which hosts the vacuum transition between two WR10 waveguides. The waveguides are connected with UG-387/UM connectors from the in- and outside of the flange. We seal the spark-eroded rectangular waveguide hole in the metal plate with a µm thick polyimide foil and an o-ring, see inset in Fig. 1 for a schematic. The helium leak rate of the box was measured to be mbarl/s, confirming the reliability of the polyimide/o-ring seal. A similar approach was proposed previously by [5]. Transmission measurements reveal an average attenuation of -4 dB and reflection below -10 dB over the full spectrum, see Fig. 3. While these values are sufficient for the planned measurements, further optimizations towards increasing the transmission, e.g. by using a thinner polyimide or mica foil, are possible. Also, no special attention was devoted to the suppression of standing waves in the transition plate, see e.g. [5] for a detailed discussion on this matter.
@@ II.2 Dielectric waveguides
DWGs, like optical fibers, rely on the principle of total internal reflection which emerges when the carrying medium (the DWG) has a substantially different dielectric constant (or refraction index) than its environment. In our approach, we use a rectangular strip of high density polyethylene (HDPE) with a dielectric constant [6] embedded in a round polyethylene (PE) foam tube with a diameter of 14 mm (). We have chosen HDPE because of its reported low loss [7], flexibility at room temperature, low piezoelectricity and low temperature dependence of the dielectric constant. In addition, HDPE material shows a good inertness to many chemicals and low moisture sensitivity. The DWG are fabricated from black or transparent HDPE sheets that are cut to a width of 2.5 mm and then rolled to a thickness of mm using an in-house made cutting tool, see Fig. A.1. The electromagnetic properties of the black and transparent DWGs are discussed below.
At mm-wave frequencies, the electromagnetic field of the signal is almost completely confined to the DWG, only a small evanescent field near the DWG’s outer boundary is present. According to simulations [2], the evanescent electric field amplitude drops of exponentially with about 6 dB/mm for a material with a comparable . DWG materials with a higher dielectric constants, like e.g. some glass reinforced plastics, could be beneficial in terms of a stronger confinement of the electric field. Considering a DWG made from a material with an , the surface electric field should drop of by 23 dB/mm [2].
To suppress crosstalk between the lines guiding the signal to and from the low-temperature stage, we place an additional metal shield around the PE foam, see Fig. 2. This is realized by a metal mesh to retain the flexible character of the DWG. The shield is especially mandatory at a high signal attenuation ( dB) sections, since without the shield the higher crosstalk leads to significant signal distortion. A suitable material for the mesh shield is stainless steal, due to its low cryogenic heat conductance, see discussion below. We measure a cross-talk suppression exceeding 60 dB in the proposed configuration. Other material solutions for the shield, like stainless steel belly-tubes or a conductive metalization on the outer side of the PE foam are also possible, but not considered here.
@@ II.4 Transmission properties of dielectric waveguides at cryogenic temperatures
For the cryogenic characterization, we employ 50 cm (25 cm) long transparent (black) DWG samples with shielding and UG-387/UM adapters on both sides. Room temperature values are compared with measurements at 2.5 K. The setup was calibrated in several consecutive cool down steps, with and without DWG samples. The data obtained by this procedure are used to de-embed the cable. Figure 2 shows the resulting transmission data. The used calibration setting is, however, sensitive to non-stationary influences, e.g. small changes of the DWG-to-adapter impedance matching. We attribute the residual oscillations with an amplitude periodicity of about 10 GHz to such an origin. Further, noise originating from the pulse tube pre-cooler is observed when measuring the high-Q resonators, see below. This noise is not present when the pre-cooler is switched off, so we suspect that the noise is caused by either vibrational or electrical interference.
@@
The following conclusions can be drawn from above findings. First, both transparent HDPE and black HDPE materials are suitable for low temperature applications. For the transparent DWG, we find a very low cryogenic attenuation, which facilitates measurements of small signals, present for instance in quantum technologies. The stronger damping of the black DWG can be used to attenuate applied mm-wave signals, suppress the background thermal radiation from the high-temperature stages, and thus enhance the noise background at the cold sample under investigation. In addition, the black HDPE filling also reduces the influence of infrared radiation, originating from warmer stages of the cryostat.
@@ II.6 Signal conditioning
To characterize the attenuation, we have measured the transmission (S21) of a transparent DWG with copper coating applied over a length of up to 65 mm in several steps. These measurements were conducted at room temperature. We observed that the attenuation is frequency independent and yields the loss of about 1 dB/mm in the range 85 - 110 GHz. The reflected signal S11 remained below -20 dB over the whole frequency range. From the low level of reflection, we conclude that the signal is mostly dissipated by the coupling of the evanescent wave to the copper powder coating. The copper coating could also be beneficial in other applications, e.g. to suppress standing waves in the cable caused, e.g., by imperfect impedance matching at the ends.
@@ III.1 Design
determines the stability of a standing wave between the mirrors. Here, the distance between the mirrors is given by and the radii of curvatures are and . Following this, one expects resonance frequencies for two curved mirrors with the same curvature when
@@
where is the speed of light and , and are the mode indices. The terms with the and functions are corrections due to the specific geometry of the cavity. The formulas above are related to the unperturbed cavity resonances, that is with no external signal coupling or mirror support structures. The coupling scheme, however, can have a strong influence on the individual modes and causes shifts of the observed resonance frequencies. One scheme often employed in the strong coupling regime makes use of a small hole drilled through one of the mirrors and to measure the reflected signal. Since we are interested in coupling quality factors above , we decided for a different approach, as depicted in Fig. 6. Here, a WR10 waveguide is passing the cavity body next to the mirrors, with a hole connecting to the cavity. The diameter of the hole is smaller than the corresponding mode wavelength in the cavity. Assuming the hole to be a circular waveguide with a cutoff frequency GHz [19], an evanescent wave can enter and leave the cavity with the signal fraction and the attenuation
@@
For the fabrication process of the niobium coated mirrors, we have taken the work [16] as a reference. First, the mirrors are machined from an oxygen-free copper block. After careful cleaning, they are annealed at °C for min in vacuum. Then, the mirrors are polished in several iterations using several diamond based polishing pastes and stainless steel matrix/pad with the inverse shape. We expect a residual surface roughness in the low m range. Residuals of the polishing paste are removed with acetone and isopropanol. Finally, a m thick coating of niobium is deposited in an argon atmosphere by DC-magnetron sputtering under rotation. Without breaking the vacuum, we add a 5 nm capping of aluminum on top of the niobium coating to prevent oxidation of the niobium.
@@ III.2 Characterization of the Fabry-Pérot cavity
Some modes with stronger coupling are already visible at room temperature and can be monitored during cool down and warm up. The frequency dependence of such a mode at around 112 GHz is shown in Fig. 7. We understand this behavior as an isometric thermal expansion of the mirror bodies and the support structure, and therefore a changing distance between the mirror surfaces. This leads to a strong temperature dependence with a shift of MHz between 40 K and room temperature. Figure 7 also shows a comparison with tabulated values for the expansion coefficient of copper. At lower temperatures, the frequency dependence flattens out (MHz shift), as expected.
At temperatures below 8 K, the superconducting state of the niobium coating (T K) becomes visible. A multitude of new modes appears with coupling quality factors ranging between and .
@@
Figure 9 shows the temperature dependence of a resonance curve in the range from 2.5 K to 6.6 K, above this temperature the resonance becomes indistinguishable from noise. With decreasing temperature the resonance frequency shifts to higher values, which we attribute to the appearance of the superconducting energy gap. We observe a shift of the resonance dip of about 500 kHz at 6.6 K, which allows us to estimate a combined superconducting kinetic inductance fraction and residual thermal expansion of the cavity to be in the order of . The amount of thermally activated quasi-particles is reduced with temperature decreasing, as indicated by the steep increase in the internal quality factor, shown in the inset.
At temperatures of 10 mK, we observe several FPC modes in the few photon power limit with internal quality factors of about . The coupling quality factors vary between individual modes between and .
@@ IV Conclusion
We have demonstrated the feasibility of employing dielectric waveguides for measurements at milli-Kelvin temperatures and low photon numbers. The DWGs are simple to manufacture, they are flexible and have low losses as well as a low heat conductance. We have verified options to attenuate applied signals at the cryogenic stages with the use of black DWGs or added copper powder coating. The black DWG also efficiently attenuates infrared radiation. With the presented implementation, thousands of DWGs could be inserted to off-the-shelf dilution refrigerators without exceeding the thermal budget. Additional metallic shields can be used to suppresses the electromagnetic cross-talk between signal lines, which is mandatory at measurements at low photon numbers. We verify the performance of the system by measuring a transmission through a superconducting Fabry-Pérot cavity at around 100 GHz, which shows internal quality factors greater than .

### v2

@@ Abstract
We propose and test a cryogenic setup comprising dielectric waveguides for mm-wave frequencies in the range of GHz and temperatures down to 10 mK. The targeted applications are quantum technologies at millimeter-wave frequencies, which require measurements at low photon numbers and noise. We show that the high density polyethylene waveguides combine a frequency independent low photon loss with a very low heat conductance. Black high density polyethylene shows a higher attenuation, which is useful to block thermal photons in a cryogenic environment. The dielectric waveguides are thermally anchored and attenuated at several stages of the cryostat. They are individually protected by additional metallic shields to suppress mutual cross-talk and external interference. With this setup, multiple superconducting resonances of a Fabry-Pérot cavity were measured at 10 mK. We find quality factors up to 15 million in the single photon limit for resonances above 100 GHz. These results show no evident influence of atomic two-level systems in the cavity.
@@ I Introduction
Millimeter-wave measurements at ultra low temperatures are widely unexplored due to several technical difficulties including, for instance, a challenging signal path thermalization and isolation. The tremendous progress in quantum computing at microwave frequencies seen over the last two decades did not so far approach the mm-wave range. The mm-wave quantum measurements would profit from an order of magnitude larger photon energy, wider bandwidth, enhanced resilience against thermal fluctuations, and may speed up qubit manipulation.
The traditionally used metallic hollow waveguides are very good conductors of mm-wave signals but are not flexible and cause an intolerable heat load on the cryogenic temperature stages. Also, the transmission of thermal radiation from elevated temperature stages may be increased due to excess flux of infrared photons through the waveguide, heating up the low-temperature stages. Here, we propose an alternative approach to 100 GHz range quantum measurements by using dielectric waveguides (DWGs) to transmit mm-wave signals from room to milli-Kelvin temperatures and back. DWGs offer several advantages: An ultra low heat conductance, very good transmission properties, and can be easily thermalized. Furthermore, DWGs exhibit very good mechanical flexibility and offer several options to handle the attenuation and suppression of infrared radiation. Although it was long known that a dielectric strip can carry mm-waves [1], practical investigations motivated by advances in information technology were carried out only about one decade ago [2] and are still an attractive area of developments [3]. Very recent usage of DWGs for cryogenic applications employed frequencies up to 300 GHz [4]. There, a circuit board was used to bridge the temperature stages of a cryostat to a base temperature of 4.2 K.
In this paper, we describe a cryogenic setup including DWG signal lines with high signal attenuation at the input and low signal attenuation at the output lines. We also report on testing the setup with a superconducting Fabry-Pérot interferometer and observation of high-Q resonances in the W-band (75-110 GHz).
@@ II Ultra-low temperature millimeter-wave setup
The developed cryogenic mm-wave setup consists of four main parts: A commercial pulse tube pre-cooled dilution refrigerator (Blufors LD250) with a base temperature of 10 mK, a mm-wave vacuum-transition, black and transparent DWGs bridging the temperature stages and a low noise amplifier with an isolator located at the 2.7 K stage. Pre-characterization measurements are also carried out in a home-made dry refrigerator with a base temperature of 2.4 K (TransMIT 2-stage pulse-tube type PtQUBE). The Fabry-Pérot cavity, discussed in the second part of the paper, is directly connected to the DWGs, see Fig. 1 for a schematic representation. All measurements presented are carried out at room temperature using a Rhode & Schwarz ZNA67 vector network analyzer (VNA) with external W-band extenders (ZC110).
@@ II.1 Vacuum transition
We use a standard KF40 flange on the top of the fridge to connect a vacuum-tight box equipped with four WR10 vacuum transitions. The stainless steel front plate of the box, where the waveguides are connected from the in- and outside using UG-387/UM connectors, is sealed using an o-ring. On the inside of the box, we use bend copper waveguides. We, further, seal each of the spark-eroded rectangular waveguide holes in the metal plate with a µm thick polyimide foil and an o-ring, see inset in Fig. 1 for a schematic. The helium leak rate of the box was measured to be mbarl/s, confirming the reliability of the polyimide/o-ring seal. A similar approach was previously proposed by [5]. Transmission measurements reveal an average attenuation of -4 dB and reflection below -10 dB over the full spectrum, see Fig. 3. While these values are sufficient for the planned measurements, further optimizations towards increasing the transmission, e.g. by using a thinner polyimide or mica foil, are possible. Also, no special attention was devoted to the suppression of standing waves in the transition plate, see e.g. [5] for a detailed discussion on this matter.
@@ II.2 Dielectric waveguides
DWGs, like optical fibers, rely on the principle of total internal reflection which emerges when the carrying medium (the DWG) has a dielectric constant (or refraction index) substantially different from its environment. In our approach, we use a rectangular strip of high density polyethylene (HDPE) with a dielectric constant [6] embedded in a round polyurethane (PU) foam tube with a diameter of 14 mm (). We have chosen HDPE because of its reported low loss [7], flexibility at room temperature, low piezoelectricity and low temperature dependence of the dielectric constant. In addition, HDPE material shows good inertness to many chemicals and low moisture sensitivity. The DWGs are fabricated from black or transparent HDPE sheets that are cut to a width of 2.5 mm and then rolled to a thickness of mm using an in-house made cutting tool, see Fig. A.1. The electromagnetic properties of the black and transparent DWGs are discussed below.
At mm-wave frequencies, the electromagnetic field of the signal is almost completely confined to the DWG, only a small evanescent field near the DWG’s outer boundary is present. According to simulations [2], the evanescent electric field amplitude drops of exponentially with about 6 dB/mm for a material with a comparable . DWG materials with higher dielectric constants, like e.g. some glass-reinforced plastics, could be beneficial in terms of a stronger confinement of the electric field. Considering a DWG made from a material with an , the surface electric field should decrease by 23 dB/mm [2].
To suppress crosstalk between the lines guiding the signal to and from the low-temperature stage, we place an additional metal shield around the PU foam, see Fig. 2. This is realized by a metal mesh to retain the flexible character of the DWG. The shield is especially mandatory at high signal attenuation sections ( dB), where the higher crosstalk would lead to significant signal distortion without the shielding. A suitable material for the mesh shield is stainless steal, due to its low cryogenic heat conductance, see discussion below. We measure a cross-talk suppression exceeding 60 dB in the proposed configuration. Other material solutions for the shield, like stainless steel corrugated tubes or a conductive metallization on the outer side of the PU foam are also possible, but not considered here.
@@ II.4 Transmission properties of dielectric waveguides at cryogenic temperatures
For the cryogenic characterization, we employ 50 cm (25 cm) long transparent (black) DWG samples with shielding and UG-387/UM adapters on both sides. Room temperature values are compared with measurements at 2.5 K. The setup was calibrated in several consecutive cool down steps, with and without DWG samples. The data obtained by this procedure is used to de-embed the cable. Figure 2 shows the resulting transmission data. The used calibration setting is, however, sensitive to non-stationary influences, e.g. small changes of the DWG-to-adapter impedance matching. We attribute the residual oscillations with an amplitude periodicity of about 10 GHz to such an origin. Further, noise originating from the pulse tube pre-cooler is observed when measuring the high-Q resonators, see below. This noise is not present when the pre-cooler is switched off, so we suspect that the noise is caused by either vibrational or electrical interference.
@@
The following conclusions can be drawn from above findings. First, both transparent HDPE and black HDPE materials are suitable for low temperature applications. Second, for the transparent DWG, we find a very low cryogenic attenuation, which facilitates measurements of small signals, present for instance in quantum technologies. Third, the stronger damping of the black DWG can be used to attenuate applied mm-wave signals, suppress the background thermal radiation from the high-temperature stages, and thus enhance the noise background at the cold sample under investigation. In addition, the black HDPE filling also reduces the influence of infrared radiation, originating from warmer stages of the cryostat.
@@ II.6 Signal conditioning
To characterize the attenuation, we have measured the transmission (S21) of a transparent DWG with copper coating applied over a length of up to 65 mm in several steps. These measurements were conducted at room temperature. We observed that the attenuation is frequency independent and yields a loss of about 1 dB/mm in the range 85 - 110 GHz. The reflected signal S11 remained below -20 dB over the whole frequency range. From the low level of reflection, we conclude that the signal is mostly dissipated by the coupling of the evanescent wave to the copper powder coating. The copper coating could also be beneficial in other applications, e.g. to suppress standing waves in the cable that may be caused by imperfect impedance matching at the ends.
@@ III.1 Design
determines the stability of a standing wave between the mirrors. Here, the distance between the mirrors is given by while and represent the radii of the curvatures. Following this, one expects resonance frequencies for two curved mirrors with the same curvature when
@@
where is the speed of light, and , and are the mode indices. The terms with the and functions are corrections due to the specific geometry of the cavity. The formulas above are related to the unperturbed cavity resonances, i.e. with no external signal coupling or mirror support structures. The coupling scheme, however, can have a strong influence on the individual modes and causes shifts of the observed resonance frequencies. One scheme often employed in the strong coupling regime makes use of a small hole drilled through one of the mirrors to measure the reflected signal. Since we are interested in coupling quality factors above , we decided for a different approach, as depicted in Fig. 6. Here, a WR10 waveguide is passing the cavity body next to the mirrors, with a hole connecting to the cavity. The diameter of the hole is smaller than the corresponding mode wavelength in the cavity. Assuming the hole to be a circular waveguide with a cutoff frequency GHz [19], an evanescent wave can enter and leave the cavity with the signal fraction and the attenuation
@@
For the fabrication process of the niobium coated mirrors, we have taken the work [16] as a reference. First, the mirrors are machined from an oxygen-free copper block. After careful cleaning, they are annealed at °C for min in vacuum. Then, the mirrors are polished in several iterations using several diamond based polishing pastes and a stainless steel matrix/pad with the inverse shape. We expect a residual surface roughness in the low µm range. Residuals of the polishing paste are removed with acetone and isopropanol. Finally, a 1 µm thick coating of niobium is deposited in an argon atmosphere by DC-magnetron sputtering under rotation. Without breaking the vacuum, we add a 5 nm capping of aluminum on top of the niobium coating to prevent oxidation of the niobium.
@@ III.2 Characterization of the Fabry-Pérot cavity
Some modes with stronger coupling are already visible at room temperature and can be monitored during cool down and warm up. The shift in center frequency (here around 112 GHz) with temperature of such a mode is shown in Fig. 7. We understand this behavior as an isometric thermal expansion of the mirror bodies and the support structure, and therefore a changing distance between the mirror surfaces. This leads to a strong temperature dependence with a shift of MHz between 40 K and room temperature. Figure 7 also shows a comparison with tabulated values for the expansion coefficient of copper. At lower temperatures, the frequency dependence flattens out (MHz shift), as expected.
At temperatures below 8 K, the superconducting state of the niobium coating (T K) becomes visible. A multitude of new modes appears with coupling quality factors ranging from to .
@@
Figure 9 shows the temperature dependence of a resonance curve in the range from 2.5 K to 6.6 K, above this temperature the resonance becomes indistinguishable from noise. With decreasing temperature the resonance frequency shifts to higher values, which we attribute to the appearance of the superconducting energy gap. We observe a shift of the resonance dip of about 500 kHz at 6.6 K, which allows us to estimate the combined superconducting kinetic inductance fraction and residual thermal expansion of the cavity to be in the order of . The corresponding change of the quality factor is shown in Fig. 10 (labeled mode B).
Mounted in the dilution refrigerator setup with a base temperature of 10 mK, we observe additional modes with higher quality factors. We have investigated two modes (”A” at 107.651 GHz and ”B” at 115.178 GHz) in detail. The two modes couple to the waveguide sufficiently weak to extract the internal quality factor from the circle fit algorithm. Their corresponding coupling quality factors are A and B.
The power dependence of mode A over a range of 50 dB, corresponding to a number of photons of about 3 to , is shown in the upper part of Fig. 10. While the internal quality factor decreases slightly from 15 M to 11 M for photon numbers below 100, the measurement averaging time in this range reaches several hours per data point (photon number), leading to larger systematic errors due to parameter drift. Over the whole range, however, the resonance frequencies lie within an interval of 7 kHz with a standard deviation of 1.7 kHz, showing no significant shift.
From these findings, we conclude that there is no apparent evidence for an influence of atomic two-level systems (TLS) on the internal quality factor. For larger photon numbers , one expects a significant increase of the internal quality factor in TLS limited resonators [23]. We attribute the absence of a TLS influence to the large mode volume and thus low electric field density at the surface of the superconducting mirrors of the FPC.
The temperature dependence of mode A and B is shown in the lower part of Fig. 10. Measured at an average number of photons, the quality factor of mode A remains constant in the range up to about 2 K, and decreases to about 1 M at 4 K. Mode B shows a similar dependence with an overall lower quality factor. We attribute this behavior to the presence of thermally activated quasi-particles, which finally suppress the modes at 7-8 K.
@@ IV Conclusion
We have demonstrated the feasibility of employing dielectric waveguides for measurements at milli-Kelvin temperatures and low photon numbers. The DWGs are simple to manufacture, they are flexible and have low losses as well as a low heat conductance. We have verified options to attenuate applied signals at the cryogenic stages with the use of black DWGs or added copper powder coating. The black DWG also efficiently attenuates infrared radiation. With the presented implementation, thousands of DWGs could be inserted to off-the-shelf dilution refrigerators without exceeding the thermal budget. Additional metallic shields can be used to suppress the electromagnetic cross-talk between signal lines, which is mandatory for measurements at low photon numbers. We verify the performance of the system by characterizing a superconducting Fabry-Pérot cavity at around 100 GHz, which shows internal quality factors greater than 15 M and coupling quality factors above 45 M down to the single photon limit.